You invested 3000 in an account that pays simple interest of 4% for 30 years.
The amount of money you'll have at the end of 30 years is
1
Expert's answer
2019-10-21T08:46:36-0400

Amount of Money Using Simple Interest


We need to find the Amount of Money at the end of 30 years using Simple Interest


Solution:

We know the formula to find the Amount using simple interest,


Amount = A = P ( 1 + rt)


Here, P = Principal Amount

r = Rate of Interest

t = time Period in Years

From the given problem,


P = 3000 , r = 4% = 0.04 , t = 30

Plug these values in the equation A  =  P ( 1 + rt)


A = 3000 ( 1 + 0.04 x 30 )


= 3000 ( 1+ 1.2) = 3000 (2.2) = 6600


Answer:

Amount of Money at the end of 30 years = 6600